Find the perimeter of a square whose area is 36.
Answer by anantha(86)

sol:
let the side be x
area of the square=side*side=x*x=x^2
according to the problem area of the square=36
x^2=36
x=sqrt(36)=6
perimeter of the square=4*side=4*6=24